OpenStudy (anonymous):

Find the interest earned in an account with $600 invested at 4.5% annual simple interest for two years.

OpenStudy (anonymous):

\[Simple Interest = \frac{Principal\times Rate\times time}{100}\] Since the rate of interest is given annually here, just put 2 (for two years) for time in the formula.

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

You are welcome :)